OpenStudy (anonymous):

Ball is kicked from a height of 24 feet at an initial velocity of 16 ft/sec. Approximately how long does it take before the ball lands on the ground?

OpenStudy (anonymous):

So far ive gotten to 0=4t^2 + 4t + but i dunno if it can factor out. Did I screw up somewhere?

OpenStudy (whpalmer4):

Unspecified is the direction in which the ball is kicked, but that is vital to answering the question...

OpenStudy (anonymous):

What??

OpenStudy (whpalmer4):

Without knowing the direction the ball gets that velocity of 16 ft/s, it is impossible to solve the problem.

OpenStudy (anonymous):

If you use the Projectile Motion Equation then it is just a matter of plugging in numbers

OpenStudy (whpalmer4):

The answer is very different if the ball goes straight up, or if it is kicked horizontally, or down toward the ground. No, it's not, unless you know the direction.

OpenStudy (anonymous):

Ok, then lets just say it is kicked upwards, now what?

OpenStudy (anonymous):

h(t) = -1/2gt^2 + vt + h

OpenStudy (whpalmer4):

Ball going straight up has equation of motion \[h(t) = -16t^2 + 16t +24\] Ball kicked horizontally from a 24 foot high platform would have equation of motion \[h(t) = -16t^2 + 0t + 24\] Okay, we'll do straight up. You're looking to find the time at which it hits the ground, or \(h(t) = 0\), so \[-16t^2 + 16t + 24 = 0\]

OpenStudy (whpalmer4):

Factor out a 4: \[-4t^2 + 4t + 6 = 0\]

OpenStudy (anonymous):

Ok, I just made a mistake in my math, I did same thing but I got -3t^2 + 4t + 8 I just messed up the 24/4

OpenStudy (whpalmer4):

I'd probably just do the quadratic formula at this point, with \(a = -4, b = 4, c = 6\) \[t = \frac{ -b \pm \sqrt{b^2 -4ac}}{2a}\]

OpenStudy (whpalmer4):

one of the values will be negative, and we can ignore that one.

OpenStudy (anonymous):

So t= 84.667?

OpenStudy (whpalmer4):

no...want to show me your work? that's off by about a factor of 45 or so

OpenStudy (whpalmer4):

that'd be amazing hang time, ball would look like it just hung there on a string or something :-)

OpenStudy (anonymous):

Calculator is really not working today :( After recalculating I got 13.166

OpenStudy (whpalmer4):

write out your numbers for me

OpenStudy (whpalmer4):

what does the equation look like before you punch it into the calculator?

OpenStudy (anonymous):

Cant write it out because of square root symbol. Hold on one sec, I will try it again

OpenStudy (whpalmer4):

just say sqrt(stuff) where you want the square root -b +/- sqrt(b^2 - 4ac) etc

OpenStudy (anonymous):

10.5? my last step was -84/-8

OpenStudy (whpalmer4):

no. I have to go now, so I'll just show you how I do it: \[a = -4, b = 4, c = 6\]\[t = \frac{ -b \pm \sqrt{b^2 -4ac}}{2a}\] \[t = \frac{-4 \pm\sqrt{(4)^2 - 4(-4)(6)}}{2(-4)} = \frac{-4\pm\sqrt{16+96}}{-8}\] you should be able to finish from there

OpenStudy (anonymous):

15

OpenStudy (anonymous):

lolol omg I hate math sooooo much thx for your help

OpenStudy (whpalmer4):

no, how do you get that? 16+96 = 112. 112 = 7*16 so sqrt(112) = sqrt(7)*sqrt(16) = 4sqrt(7) answer is \[\frac{-4\pm4\sqrt{7}}{-8}\]

OpenStudy (anonymous):

Ohhh yea the sgrt. I forgot about it

OpenStudy (anonymous):

ok, that makes a lot more sense

OpenStudy (anonymous):

Would that also be the maximum height that the ball can reach?

OpenStudy (whpalmer4):

No, the maximum height will be at the vertex of the parabola. If you have a parabola written in the form \[y = ax^2 + bx + c\]the vertex x-coordinate is given by \[x = -\frac{b}{2a}\]and the vertex y-coordinate can be found by plugging that value of \(x\) back into the equation. Also, the vertex is exactly at the midpoint on the x-axis between the two solutions where the parabola = 0. Note that this is NOT necessarily half of the flight time! If you start at a different elevation than you land, the vertex will be before the halfway mark in the flight.